OpenStudy (anonymous):

Part 1: Find the solution to the equation log16 32 = x + 3. (4 points) Part 2: Using complete sentences, explain the procedure used to solve this equation. (4 points)

OpenStudy (campbell_st):

ok... you need to use the log law which is \[a^{\log{a}(b)} = b\] so if you raise each side as a power of 16 you will have \[32 = 16^{x + 3}\] next you need to write 32 and 16 in terms of powers of 2 so you have \[2^5 = (2^4)^{x + 3}\] use the power of a power and you get \[2^5 = 2^{4x + 12}\] so all you need to do is solve 5 = 4x + 12 to find the value of x